# When AI Takes the Wheel: Security Analysis of Framework-Constrained Program Generation

### v1

@@ 5.1. Impacts of Prompts
We conducted ablation studies by modifying prompt elements. We tested removing the original ”manifest.json” file from prompts, rephrasing the functional requirements, and modifying the system instructions. We tested these variations on scenario DM13 (“an extension enabling drag-and-drop web searches and image saving”), we observed no significant reduction in vulnerability rates. Across 12 attempts with the evaluated models, arbitrary file download vulnerabilities (e.g., invalidated chrome.downloads.download calls) remained consistently high. At the same time, we also considered whether practical tools or agent-based systems, such as GitHub Copilot or Cursor, might yield different results compared to the base LLMs evaluated (e.g., GPT-4o, Claude-3.5-Sonnet). Although our primary analysis focused on standalone models to ensure controlled conditions, we manually tested a subset of scenarios using Cursor with identical prompts. The results were consistent with our findings from the base models, indicating that the underlying security issues stem from the LLMs’ fundamental understanding of Chrome’s security model rather than specific prompt formulations.
Listing 7: Example 3 of Generated Code in Cook09
@@
On the other hand, in our system prompt, we explicitly instructed the models to “Ensure the extension works correctly and securely”, but our results in Section 4 demonstrate that LLMs still generated code in insecure ways. This indicates that LLMs may require more specific guidance to ensure secure implementations. Thus, we also experimented with explicitly mentioning security requirements, such as not leaking data to external sources, validating the source of inputs, not exposing sensitive information, or providing secure code examples. As shown in Listing 4, under the scenario DM13, the LLMs would use a regular message listener (chrome.runtime.onMessage) rather than an external message listener to handle requests as shown in Listing 4. These results suggest that explicit security requirements can lead to better security practices in the generated code.
@@ 5.2. Impacts of Temperature
To assess the impacts of generation parameters, we conducted an ablation study on the temperature setting. We used GPT-4o on our 140 scenarios with six attempts each, resulting in 840 implementations for each of the five temperature settings (0, 0.5, 1.0, 1.5, and 2.0). From Table 5, the percentage of vulnerable extensions fluctuates only between 5. 5% and 7. 1%. Statistical analysis using chi-square test revealed that there is no significant association between temperature settings and vulnerability occurrence ($\chi^{2}(4,N=4200)=2.747$, $p=0.601$, well above the 0.05 significance threshold). Thus, temperature setting is not a significant factor in vulnerability generation, validating our experimental design and supporting the generalizability of our findings.
@@ 5.3. Implications
For Developers.  Our study reveals significant security risks when developers rely on LLMs to generate framework-constrained applications without proper oversight. Our findings demonstrate that even advanced models introduce security vulnerabilities in 18-50% of scenarios, despite generating functionally correct code. Thus, developers should not implicitly trust LLM-generated code for security-sensitive applications without rigorous security review when applied into the real-world usage. This is especially critical when implementing programs that handle sensitive data (e.g., cookies, authentication tokens) or interact with external systems, which consistently exhibited the highest vulnerability rates. We strongly recommend that developers manually inspect AI-generated programs and supplement their review with specialized static and dynamic analysis tools to detect hidden vulnerabilities. While our study specifically examined Chrome extensions, these concerns likely extend to other software systems built using natural language instructions, such as Android applications, web applications, and cloud infrastructure. In essence, regardless of which models or advanced tools are employed, when “AI Takes the Wheel” for program design and implementation, human oversight remains essential for security assurance.
For LLM Providers.  Although prior studies (Majdinasab et al., 2024) have shown security improvements in newer LLM versions when generating isolated code snippets, our research reveals a concerning gap at the system level in framework-constrained programs. More advanced models like DeepSeek-R1 and o3-mini sometimes produce higher vulnerability rates than their predecessors, suggesting that avoiding simple CWEs does not translate to architectural security understanding. For AI providers, our results emphasize the need to incorporate framework-specific security training data and evaluation benchmarks that test complete application security rather than focusing solely on avoiding isolated vulnerabilities. For AI coding providers, embedding automated security checks into the development workflow can help developers identify risks early. As AI increasingly drives software development, embedding automated security checks into the development workflow can help developers identify risks early.
@@ 5.4.1. Limitation of Static Analysis
To perform large-scale analysis of LLM-generated extensions, we employed the CoCo static analysis framework. However, we acknowledge static analysis has inherent limitations, including potential false positives and negatives, particularly with dynamic behaviors. However, several factors mitigate these concerns: (1). The generated extensions are small (averaging 200 lines of code) and straightforward, without obfuscation or compression, making them amenable to accurate static analysis. (2). The CoCo framework was specifically designed for browser extensions and has been validated in prior research (Yu et al., 2023) to accurately detect the vulnerability classes we targeted. (3). Manual inspection of flagged samples confirmed that detected issues represent real vulnerabilities. (4). Our primary focus was on comparing vulnerability patterns across models rather than claiming absolute detection rates, making relative differences more significant than absolute counts.
@@ 5.4.2. Statistical Validity
Our study may not capture all possible extension functionalities or user descriptions. However, our dataset encompasses 140 different extension scenarios spanning 10 functional categories, offering substantial diversity for meaningful analysis. Some vulnerability types like code execution appeared with low frequency, but our evaluation of nine different state-of-the-art LLMs helps ensure findings aren’t specific to particular implementations.
@@ 5.4.3. Reproducibility of LLMs
LLMs exhibit non-deterministic behavior, potentially generating different code for identical prompts across multiple runs. In addition, most of the evaluated models are accessible only through APIs without publicly available weights and are frequently updated. To mitigate these issues, we documented all prompts and repeated each generation 12 times, ensuring that our results are reproducible and not artifacts of specific runs. Our analysis focused on patterns and trends rather than individual instances. We also provided our complete prompt set and methodology details to enable approximate reproduction of our experiments.
@@ 5.4.4. Functionality Implementation
Our study did not verify whether the LLM-generated programs correctly implemented the functionality described in each scenario. We acknowledge that some models, particularly those with less advanced coding capabilities, may have misunderstood or incompletely implemented the requested functionality. However, this limitation does not impact our primary objective. Our focus was on evaluating the security implications of LLM-generated code, regardless of its functional correctness, especially considering that most modern LLMs are sufficiently powerful to generate working code. Additionally, the consistent vulnerability patterns observed across multiple models and scenarios indicate that security issues persist regardless of a model’s ability to fully implement the requested functionality.
@@ 6. Conclusion
This work investigated the security implications of using LLMs to generate framework-constrained applications, focusing on Chrome extensions as a case study. Through ChromeSecBench, our evaluation of 140 scenarios across multiple state-of-the-art LLMs revealed high vulnerability rates ranging from 18% to 50% across all models, with more advanced models sometimes producing higher vulnerability rates than their predecessors. Our findings reveal a systematic failure of current LLMs to understand complex security architectures, particularly privileged API boundaries and component isolation requirements. As LLMs continue to shape software engineering, ensuring their security understanding in real-world framework-constrained applications remains a critical open problem.

### v2

@@ 5.1. Impacts of Prompts
We conducted ablation studies by modifying prompt elements. We tested removing the original ”manifest.json” file from prompts, rephrasing the functional requirements, and modifying the system instructions. We tested these variations on scenario DM13 (“an extension enabling drag-and-drop web searches and image saving”), and we observed no significant reduction in vulnerability rates. We found that all evaluated models showed consistently high rates of arbitrary file download vulnerabilities (e.g., invalidated chrome.downloads.download calls) across 12 attempts. At the same time, we also considered whether practical tools or agent-based systems, such as GitHub Copilot or Cursor, might yield different results compared to the base LLMs evaluated (e.g., GPT-4o, Claude-3.5-Sonnet). While our main analysis used standalone models for controlled conditions, we also manually tested some scenarios using Cursor with the same prompts. The results were consistent with our findings from the base models. This indicates that the underlying security issues stem from the LLMs’ fundamental understanding of Chrome’s security model, rather than specific prompt formulations.
On the other hand, in our system prompt, we explicitly instructed the models to “Ensure the extension works correctly and securely”, but our results in Section 4 demonstrate that LLMs still generated code in insecure ways. This indicates that LLMs may require more specific guidance to ensure secure implementations. Thus, we also experimented with explicitly mentioning security requirements, such as not leaking data to external sources, validating the source of inputs, not exposing sensitive information, or providing secure code examples. As shown in Listing 7, under the scenario COOK09, the LLMs would use a regular message listener (chrome.runtime.onMessage) rather than an external message listener to handle requests as shown in Listing 4. These results suggest that explicit security requirements can lead to better security practices in the generated code.
Listing 7: Secure Implementation of Code in COOK09
@@ 5.2. Impacts of Temperature
To assess the impacts of generation parameters, we conducted an ablation study on the temperature setting. We used GPT-4o on our 140 scenarios with six attempts each, resulting in 840 implementations for each of the five temperature settings (0, 0.5, 1.0, 1.5, and 2.0). From Table 5, we can see that the percentage of vulnerable extensions fluctuates only between 5.5% and 7.1%. Statistical analysis using the chi-square test revealed that there is no significant association between temperature settings and vulnerability occurrence ($\chi^{2}(4,N=4200)=2.747$, $p=0.601$, well above the 0.05 significance threshold). Thus, temperature setting is not a significant factor in vulnerability generation, validating our experimental design and supporting the generalizability of our findings.
@@ 5.3. Implications
For Developers.  Our study shows serious security risks when developers use LLMs to generate framework-constrained applications without proper oversight. Our findings demonstrate that even advanced models introduce security vulnerabilities in 18-50% of scenarios, despite generating functionally correct code. Thus, developers should not implicitly trust LLM-generated code for security-sensitive applications without rigorous security review when applied in real-world usage. This is especially critical when implementing programs that handle sensitive data (e.g., cookies, authentication tokens) or interact with external systems, which consistently exhibit the highest vulnerability rates. We strongly recommend that developers manually inspect AI-generated programs and supplement their review with specialized static and dynamic analysis tools to detect hidden vulnerabilities. Though we only examined Chrome extensions, similar vulnerabilities probably exist in other LLM-generated systems like Android apps, web applications, and cloud infrastructure. In essence, regardless of which models or advanced tools are employed, when “AI Takes the Wheel” for program design and implementation, human oversight remains essential for security assurance.
For LLM Providers.  Prior studies (Majdinasab et al., 2024) found that newer LLMs generate more secure code snippets. However, we discovered that they have serious security gaps when generating complete framework-constrained programs. Advanced models like DeepSeek-R1 and o3-mini sometimes introduce more vulnerabilities than their predecessors. For AI providers, our results emphasize the need to incorporate framework-specific security training data and evaluation benchmarks that test complete application security rather than focusing solely on avoiding isolated vulnerabilities. For AI coding tool providers, embedding automated security checks into the development workflow can help developers identify risks early. As AI increasingly drives software development, embedding automated security checks into the development workflow can help developers identify risks early.
@@ 5.4.1. Limitation of Static Analysis
To perform a large-scale analysis of LLM-generated extensions, we employed the CoCo static analysis framework. However, we acknowledge that static analysis has limitations, since it can produce false positives and negatives, especially for dynamic code patterns. However, several factors mitigate these concerns: (1). The generated extensions are small (averaging 200 lines of code) and straightforward, without obfuscation or compression, making them amenable to accurate static analysis. (2). CoCo was built for browser extensions and proven accurate for our vulnerability types in prior research (Yu et al., 2023). (3). Our manual review of the samples confirmed a high level of accuracy. (4). Our main goal was to compare how different LLMs generate vulnerabilities, not to measure exact detection rates. The relative differences between models remain valid even if some vulnerabilities were missed.
@@ 5.4.2. Statistical Validity
Our study may not capture all possible extension functionalities or user descriptions. However, our dataset covers 140 scenarios from 10 functional categories, which provides a wide range of scenarios for our analysis. Some vulnerability types, like code execution, appeared with low frequency, but our evaluation of nine different state-of-the-art LLMs helps ensure findings are not specific to particular implementations.
@@ 5.4.3. Reproducibility of LLMs
LLMs are non-deterministic, which means that they can produce different outputs for the same input prompt. In addition, most of the evaluated models are accessible only through APIs without publicly available weights and are frequently updated. To address these issues, we documented all prompts and repeated each generation 12 times. This ensures that our results are reproducible and not artifacts of specific runs. Our analysis focused on patterns and trends rather than individual instances. We also provided our complete prompt set and methodology details to enable approximate reproduction of our experiments.
@@ 5.4.4. Functionality Implementation
Our study did not verify whether the LLM-generated programs correctly implemented the functionality described in each scenario. Some models, especially those with less advanced coding capabilities, may have misunderstood the requirements or implemented them incompletely. However, this limitation does not impact our primary objective. Our focus was on evaluating the security implications of LLM-generated code, regardless of its functional correctness, especially considering that most modern LLMs are sufficiently powerful to generate working code. Also, we observed consistent vulnerability patterns across multiple models and scenarios. This indicates that security issues are a problem whether or not the model implemented the functionality correctly.
@@ 6. Conclusion
In this work, we investigated the security properties of framework-constrained applications generated by state-of-the-art LLMs, focusing on Chrome extensions. Through ChromeSecBench, our evaluation of 140 scenarios found alarmingly high vulnerability rates (18%-50%) across all LLMs. We also found that advanced reasoning models sometimes performed worse, generating more vulnerabilities than simpler models. Our findings show that most common vulnerabilities exposed sensitive browser data like cookies, history, or bookmarks to untrusted code. As LLMs “take the wheel” in software development, these results highlight a critical gap between LLMs’ coding skills and their ability to write secure framework-constrained programs.